### v1

# Improving Node Representation by Boosting Target-Aware Contrastive Loss

Abstract.
Graphs model complex relationships between entities, with nodes and edges capturing intricate connections. Node representation learning involves transforming nodes into low-dimensional embeddings. These embeddings are typically used as features for downstream tasks. Therefore, their quality has a significant impact on task performance. Existing approaches for node representation learning span (semi-)supervised, unsupervised, and self-supervised paradigms. In graph domains, (semi-)supervised learning often only optimizes models based on class labels, neglecting other abundant graph signals, which limits generalization. While self-supervised or unsupervised learning produces representations that better capture underlying graph signals, the usefulness of these captured signals for downstream target tasks can vary. To bridge this gap, we introduce Target-Aware Contrastive Learning (Target-aware CL) which aims to enhance target task performance by maximizing the mutual information between the target task and node representations with a self-supervised learning process. This is achieved through a sampling function, XGBoost Sampler (XGSampler), to sample proper positive examples for the proposed Target-Aware Contrastive Loss (XTCL). By minimizing XTCL, Target-aware CL increases the mutual information between the target task and node representations, such that model generalization is improved. Additionally, XGSampler enhances the interpretability of each signal by showing the weights for sampling the proper positive examples. We show experimentally that XTCL significantly improves the performance on two target tasks: node classification and link prediction tasks, compared to state-of-the-art models. Our code is available on GitHub [^1] Github link will be provided after publication to preserve anonymity..
Graph Neural Network, Contrastive Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; October,2024; arXiv
1. Introduction
A graph is a data structure that represents relationships between entities or elements using nodes and edges. Node representation learning involves converting nodes in a graph into lower-dimensional embeddings. Node embeddings are frequently utilized as features for downstream target tasks. Therefore, the performance of these tasks is greatly influenced by the quality of the learned embeddings. Node representation learning methods are optimized using various loss functions, including (semi-)supervised, unsupervised, or self-supervised. (Semi-)Supervised learning involves accessing ground truth information associated with a downstream task during training. However, when downstream class labels are available, the optimization may overfit to those labels and underutilize other, more abundant graph signals, which can reduce generalization performance.
Figure 1. Model space consisting of models trained by different supervision signals (text labels in each model space). The optimal model for a target task is represented by a yellow star. Different supervision signals may result in varying degrees of overfitting or generalization error. Our goal is to increase the likelihood of finding the optimal model by optimizing it through the use of labels and self-supervision in combination.
Unsupervised and self-supervised learning use the underlying patterns or graph structure for optimization. In this case, the node representations are learned to capture graph signals in training. However, the relationship between these graph signals and a downstream target task is not always clear. Optimizing for irrelevant information can negatively impact task performance (Tschannen et al., 2020), as shown in Figure 3.
While previous studies have demonstrated that self-supervised learning can greatly enhance task performance when they are relevant to the target task (Suresh et al., 2021b), there has been relatively little work focusing on analyzing the relevance between self-supervised methods and the target task. Therefore, in this work we propose a self-supervised approach, Target-Aware Contrastive Learning (Target-aware CL), which can actively adapt its contrastive loss function based on the target task. Given a pair/set of positive examples and negative examples, the general goal of contrastive loss is to bring positive examples together and negative examples apart from each other. Unlike self-supervised learning, Target-aware CL uses Target-Aware Contrastive Loss (XTCL), which involves strategically selecting the positive examples that enhance the mutual information between the target task and node representations, such that XTCL is relevant to the target task.
However, there are two main challenges for strategically selecting the positive examples that increase the mutual information between the target task and node representations. First, how each node label relates to a target task depends on its local neighborhood, e.g., a local subgraph, and on its node attributes. Instead of proposing a general positive example sampling strategy as in previous studies (Hagerup et al., 1993), one sampling strategy should be learned for each node. However, it’s inefficient to learn one strategy for each node, especially when the size of the graph is large. Second, the amount of ground truth is usually limited. It’s even more difficult to learn a sampling strategy for each node if limited ground truth is available.
To tackle these challenges, we develop an XGBoost Sampler (XGSampler) that uses limited ground truth labels in the training data to learn how to sample positive examples that increase the mutual information between the target task and node representations. Our use of XGSampler not only increases the likelihood of sampling nodes to improve the relevance between the target task and node representations, but it also enhances the interpretability of models by indicating the importance of each graph signal for the target task. Our main contributions of this work are as follows:
  • We introduce a novel concept of supervision for representation learning (Target-aware CL), which integrates target-relevant graph signals into self-supervised optimization.
  • We propose Target-Aware Contrastive Loss (XTCL) as an example of Target-aware CL that aims to make the contrastive loss more relevant to a target task. It achieves this by utilizing XGSampler to sample proper positive examples that enhance the mutual information between the target task and node representations.
  • We evaluate XTCL with GNNs empirically, and show that it can reduce generalization error and enhance performance significantly on both node classification and link prediction compared with state-of-the-art models.
2. Notations and Preliminaries
We first introduce basic concepts and define common notations in this section. A general graph is defined as ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}},{\bm{X}},{\bm{Y}})$ where ${\mathcal{V}}$ is the node set, ${\mathcal{E}}:{\mathcal{V}}\times{\mathcal{V}}$ denotes all edges, ${\bm{X}}\in{\mathbb{R}}^{|{\mathcal{V}}|\times m}$ is node attributes ($m$ is the dimension), and ${\bm{Y}}=\{y_{1},y_{2}\cdots,y_{|{\mathcal{V}}|}\}$ is the set of node labels in which class label $y_{u}\in\{1,2,\cdots,C\}$. Such graphs contain various semantic relations =${\mathcal{R}}=\left\{r_{1},r_{2},\cdots,r_{|{\mathcal{R}}|}\right\}$ which refer to various relationships between pairs of nodes, including structural similarities such as shortest path distance or attribute similarities such as neighbor label distribution. We use these semantic relation as our graph signals in this paper. For each relation $r\in{\mathcal{R}}$, we define a similarity measure $s_{r}:{\mathcal{V}}\times{\mathcal{V}}\rightarrow[0,1]$ to record the strength of the relation between two nodes and $S_{r}$ is the random variable describing the similarity values for $r$.
2.1. Node Contrastive Learning
Given a graph ${\mathcal{G}}$, a graph-based model $f:{\mathcal{G}}\rightarrow{\mathbb{R}}^{|{\mathcal{V}}|\times h}$ transforms nodes in the graph into node representation $Z=f({\mathcal{G}})$, where $h$ is the embedding dimension. We denote $z_{u}$ as embedding of node $u\in{\mathcal{V}}.$ Specifically, $Z$ is optimized by Contrastive Loss (CL) (Chuang et al., 2020; Saunshi et al., 2019; Xiong et al., 2020) $\mathcal{L}(u,v_{p},\left\{v_{n_{k}}\right\}_{k=1}^{K})$, which is generally defined as:
$\mathcal{L}(u,v_{p},\left\{v_{n_{k}}\right\}_{k=1}^{K})=-\log\frac{exp\left(z_{u}^{\top}z_{v_{p}}\right)}{exp\left(z_{u}^{\top}z_{v_{p}}\right)+\sum_{k=1}^{K}exp\left(z_{u}^{\top}z_{v_{n_{k}}}\right)},$
where $u,v_{p},v_{n_{k}}\in{\mathcal{V}}$, and $K$ is the number of negative examples We will refer to $u$ as the query node, $v_{p}$ as positive examples, and $v_{n_{k}}$ as negative examples. The positive example $v_{p}$ for $u$ is usually sampled from some probability function w.r.t $u$, e.g. $v_{p}$ sampled based on the probability that a edge exist between $u$ and $v_{p}$ (Mo et al., 2022; Kipf and Welling, 2017a) or the probability that two nodes shares similar topological structure (Dong et al., 2022), etc. Conversely, negative examples are randomly sampled from ${\mathcal{V}}$ or nodes distinct from $u$. If $B>1$ positive examples are sampled for query node $u$, the contrastive loss
$\mathcal{L}$ $\bigg{(}u,\left\{v_{p_{b}}\right\}_{b=1}^{B},\left\{v_{n_{k}}\right\}_{k=1}^{K}\bigg{)}=$
(1) $-\frac{1}{B}\sum_{b=1}^{B}\log\Bigg{[}\frac{exp\left(z_{u}^{\top}z_{v_{p_{b}}}\right)}{exp\left(z_{u}^{\top}z_{v_{p_{b}}}\right)+\sum_{k=1}^{K}exp\left(z_{u}^{\top}z_{v_{n_{k}}}\right)}\Bigg{]}$
is minimized. The contrastive loss encourages positive examples to be embedded close to $u$ and discourages negative examples from being embedded close to $u$.
Depending on the sampling strategy for positive examples, the contrastive loss can be categorized into self-supervised (Dong et al., 2022; Kipf and Welling, 2017a; Mo et al., 2022) and supervised (Khosla et al., 2020) contrastive loss.
Definition 0.
Self-Supervised Contrastive Loss (SSCL). A graph-based model $f$ is usually obtimized by SSCL with positive examples sampled based on a subset semantic relations ${\mathcal{R}}_{f}\subseteq{\mathcal{R}}.$ $v_{p}^{r}$ can be sampled from the set $\{v\>|\>s_{r}(u,v^{r}_{p})>\eta_{r}\}$ for $r$. The threshold $\eta_{r}\in(0,1]$ is used to select examples with high similarity w.r.t $r$ for $u$. If $B>|{\mathcal{R}}_{f}|>1$, previous studies usually uniformly select $\lfloor B/|{\mathcal{R}}_{f}|\rfloor$ positive samples for each $r\in{\mathcal{R}}_{f},$ e.g.
$\left\{v_{p_{b}}\right\}_{b=1}^{B}=\cup_{r\in{\mathcal{R}}_{f}}\{v_{p_{i}}^{r}\}_{i=1}^{\lfloor B/|{\mathcal{R}}_{f}|\rfloor}.$
The negative examples are sampled randomly from ${\mathcal{V}}$. Then, these positive and negative examples can be utilized to minimize the loss in equation 2.1. Depending on which relation is used, the SSCL model space is illustrated by the blue areas in Figure 1.
Let ${\mathcal{I}}(X_{1};X_{2})$ be the mutual information between two random variables $X_{1}$ and $X_{2}$. According to (van den Oord et al., 2018), minimizing the loss in equation 2.1 with the equal number of positive examples, i.e. $\lfloor B/|{\mathcal{R}}_{f}|\rfloor$, for each $r\in{\mathcal{R}}_{f}$ is equivalent to uniformly maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$.
Definition 0.
Supervised Contrastive Loss (SCL). In general SCL uses ground truth labels to sample positive examples, e.g.
$\left\{v_{p_{b}}\right\}_{b=1}^{B}\subset\{v\>|\>y_{u}=y_{v},u,v\in V_{L}\}$
where $V_{L}\subseteq{\mathcal{V}}$ is the set of labeled nodes, and the negative examples are sampled randomly from ${\mathcal{V}}$. The SCL loss can be derived from equation 2.1 by utilizing these positive and negative examples, and the models optimized by SCL is in the model space depicted by the red (training) area in Figure 1. By sampling positive examples based on node labels ${\bm{Y}}$, minimizing the loss in equation 2.1 is equivalent to maximizing ${\mathcal{I}}({\bm{Y}};Z)$ (van den Oord et al., 2018).
Node representation $Z$ learned by minimizing various loss functions can be used for downstream prediction by learning another target-specific model. A target-specific model produces outcomes for a target task by learning a mapping function between $Z$ and the target. For example, a classifier is a target-specific model mapping $Z$ to class labels for nodes.
3. Learning Node Representations with Target-Aware Contrastive Loss
In this section, we will discuss the motivation for Target-Aware Contrastive Learning (Target-aware CL) for node representations and then describe our implementation of Task-aware Contrastive Loss (XTCL) which uses XGSampler for Target-aware CL.
3.1. Target-Aware Contrastive Loss
As discussed in (van den Oord et al., 2018), minimizing the loss in equation 2.1 by sampling the equal number of positive examples for each $r\in{\mathcal{R}}_{f}$ is equivalent to uniformly maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}.$ We argue that the generalization of a graph-based model $f$ is impacted by whether uniformly maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ is equivalent to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. Our main argument is that there is a high chance that uniformly maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ is not equivalent to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. Consequently, the node representation may be considered inadequate for the intended classification task. We will empirically verify this statement in Section 5.3.
Instead, each ${\mathcal{I}}(S_{r};Z)$ for $r\in{\mathcal{R}}_{f}$ should be maximized according to some weight, which can be controlled by the number of positive nodes. A possible approach is to decide the number of positive examples for each $r$ based on the relation between $r$ and the target task. To make the relation between $r$ and the target task more clear, we define:
Definition 0.
Task Positive Function. A task positive function ${\mathbb{I}}_{t}(u,v)\in\{0,1\}$ is defined to verify whether two nodes have the same target-relevant features. If ${\mathbb{I}}_{t}(u,v)=1$, then $u$ and $v$ share the same target-relevant feature, or vice versa. If the task is node classification, ${\mathbb{I}}_{t}(u,v)={\mathbb{I}}\{y_{u}=y_{v}\}$. On the other hand, if the task is link prediction, ${\mathbb{I}}_{t}(u,v)={\mathbb{I}}\{(u,v)\in\mathbb{E}\}$.
The target task can be any type of task as long as the task positive function is defined for the task $t$.
Specifically, for a node $u$, the number of positive examples of $r$ can be determined by
(2) ${\mathbb{P}}_{u}\big{(}{\mathbb{I}}_{t}(u,v)=1\big{|}s_{r}(u,v)\big{)}.$
This probability quantifies that how likely to sample two nodes with ${\mathbb{I}}_{t}(u,v)=1$ by knowing the similarity values between $u$ and $v$. If the probability is high, a sampling strategy for positive examples should sample more nodes w.r.t $r$, otherwise sampling less nodes w.r.t $r$.
Note that equation 2 is defined w.r.t to each node. The probability in equation 2 should be different given the same similarity values because the strength of the relationship between nodes is contingent upon the local subgraph of $u$. For instance, it is possible that $s_{r}(u,v)=0.1$ informs higher probability that ${\mathbb{I}}_{t}(u,v)=1$ for $u$ when most of nodes in the subgraph having $s_{r}(u,v)=0$, while $s_{r}(u,v)=0.1$ informs lower probability that ${\mathbb{I}}_{t}(u,v)=1$ when most of nodes with similarity values greater than 0.4.
Sampling the number of positive examples based on equation 2 for each $r\in{\mathcal{R}}_{f}$ is equivalent to maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ based on weights introduced by equation 2. We hypothesis that weighted maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ based on equation 2 is similar to maximizing ${\mathcal{I}}({\bm{Y}};Z)$ and therefore model generalization is improved.
Algorithm 1 XGBoost Sampler Training
0:  Ordered semantic relation set ¯=$\overline{{\mathcal{R}}}=\{r_{1},r_{2},\cdots,r_{|{\mathcal{R}}|}\}$, training instances $U_{L}$
0:  Classifiers ${f_{u}|u\in{\mathcal{V}}}$
1:  Initialize all $\hat{y}^{(0)}_{u,v}=0$
2:  for all $i\in\{1,2,\cdots,|{\mathcal{R}}|\}$ do
3:     for all $(u,v)\in U_{L}\times U_{L}$ do
4:        $g_{u,v}\leftarrow\sigma(\hat{y}^{(i-1)}_{u,v})-y_{u,v}$
5:        $h_{u,v}\leftarrow\sigma(\hat{y}^{(i-1)}_{u,v})(1-\sigma(\hat{y}^{(i-1)}_{u,v}))$
6:     end for
[…]
9:     for all $(u,v)\in U_{L}\times U_{L}$ do
10:        if $s_{r_{i}}(u,v)>\eta_{r_{i},u}$ then
11:           $\hat{y}^{(i)}_{u,v}=\hat{y}^{(i-1)}_{u,v}+w_{r_{i},1}$
12:        else
13:           $\hat{y}^{(i)}_{u,v}=\hat{y}^{(i-1)}_{u,v}+w_{r_{i},0}$
14:        end if
15:     end for
16:  end for
Based on the above hypothesis, our Task-aware Contrastive Loss (XTCL) will actively adapt the number of positive examples based on equation 2 for each $r\in{\mathcal{R}}_{f}$ w.r.t a target task. By minimizing XTCL for $f$, it can maximize ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ based on equation 2. Therefore, $f$ can generalize better because maximize ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ based on equation 2 is similar to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. Then our Target-Aware Contrastive Learning is defined as:
Definition 0.
Target-Aware Contrastive Learning. (TCL) TCL uses ${\mathbb{P}}_{u}({\mathbb{I}}_{t}(u,v)=1|s_{r}(u,v))~{}\forall r\in{\mathcal{R}}_{f}$ to identify positive examples for equation 2.1 based on,
(3) $\left\{v_{p_{b}}\right\}_{b=1}^{B}=\mathop{\mathrm{arg\,top}\,k}_{\{v\}}\>\sum_{r\in{\mathcal{R}}_{f}}{\mathbb{P}}_{u}\bigg{(}{\mathbb{I}}_{t}(u,v)=1\bigg{|}s_{r}(u,v)\bigg{)}$
where $k=B$ and $\mathop{\mathrm{arg\,top}\,k}$ returns the $\{v\}$ that result in the top-$k$ probability values.
However, it is usually difficult to know ${\mathbb{P}}_{u}\big{(}{\mathbb{I}}_{t}(u,v)=1\big{|}s_{r}(u,v)\big{)}$ because ${\mathbb{I}}_{t}(u,v)$ for most of nodes is unknown with limited ground truth labels. Also, learning ${\mathbb{P}}_{u}\big{(}{\mathbb{I}}_{t}(u,v)=1\big{|}s_{r}(u,v)\big{)}~{}\forall r\in{\mathcal{R}}_{f}$ for each $u$ is computationally burdensome and time-consuming. To tackle this challenge, we propose XGBoost Sampler (XGSampler) to sample positive examples by equation 3 such that minimizing equation 2.1 is similar to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. As we will show later in Section 3.2, XGSampler is a more efficient approach in practice.
Since equation 3 is similar to XGBoost (Chen and Guestrin, 2016), the proposed XGSampler is an ensemble learning algorithm that combines multiple simple regression trees $f_{r,u}$ into one strong XGSampler $f_{u}$ to select positive examples for $u$. A relation regression tree is defined as:
Definition 0.
Relation Regression Tree (RelTree). For semantic relation $r\in{\mathcal{R}}$, a relation regression tree is
(4) $f_{r,u}(v)=\begin{cases}w_{r,0},&\text{ if }s_{r}(u,v)<\eta_{r,u}\\ w_{r,1},&\text{ if }s_{r}(u,v)\geq\eta_{r,u},\end{cases}$
where $w_{r,0}\in{\mathbb{R}}$ and $w_{r,1}\in{\mathbb{R}}$.
Note that $\eta_{r,u}$ in $f_{r,u}(v)$ is determined for each node depending on the local subgraph of $u$ and $\eta_{r,u}$ is determined according to the similarity values $s_{r}(u,v)$ between query node $u$ and all other nodes in the graph. In our experiments, $\eta_{r,u}$ is set as the $99$-percentile of values of $\{s_{r}(u,v)|v\in{\mathcal{V}}\}$ given query nodes $u$. While $w_{r,0}$ and $w_{r,1}$ is the same for all nodes in the regression tree of $r$, this approach increase generalization of $f_{r,u}$ without overfitting to the local subgraph of $u$. XGSampler is defined as: $f_{u}(v)=\sigma(y_{u,v}),~{}y_{u,v}=\sum_{i=1}^{|{\mathcal{R}}_{f}|}f_{r_{i},u}(v)$. Here $y_{u,v}\propto\sum_{r\in{\mathcal{R}}_{f}}{\mathbb{P}}_{u}\big{(}{\mathbb{I}}_{t}(u,v)=1\big{|}s_{r}(u,v)\big{)}$ in equation 3. The top-$B$ nodes having higher $f_{u}(v)$ are selected as the positive examples for minimizing equation 2.1.
Since XGBoost learns a regression tree iteratively, our target for each training iteration is to learn $w_{r,0}$,$w_{r,1}$ for RelTree based on the classifier from the previous iteration:
(5) $\hat{y}^{(\tau)}_{u,v}=\sum_{i=1}^{\tau-1}f_{r_{i},u}(v)+f_{r_{\tau},u}(v)=\hat{y}^{(\tau-1)}_{u,v}+f_{r_{\tau},u}(v).$
equation 5 means that the prediction of $\hat{y}^{(\tau)}_{u,v}$ is the combination of current RelTree $f_{r_{\tau},u}(v)$ and its previous RelTrees $\sum_{i=1}^{\tau-1}f_{r_{i},u}(v)$. At iteration $\tau$, RelTree $f_{r_{\tau},u}(v)$ for relation $r_{\tau}$ is learned from the regularized objective:
(6) $\mathcal{L}^{(\tau)}_{xg}=\sum_{u\in U_{L}}\Bigg{[}\sum_{v\in U_{L}}\ell\left(y_{u,v},\;\hat{y}^{(\tau-1)}_{u,v}+f_{r_{\tau},u}(v)\right)+\Omega(f_{r_{\tau},u})\Bigg{]},$
where $U_{L}$ is determined based on the target task and $y_{u,v}={\mathbb{I}}_{t}(u,v)$ and $\Omega(f_{r_{\tau},u})=\lambda(w^{2}_{r_{\tau},0}+w^{2}_{r_{\tau},1})/2$, where $\lambda\in{\mathbb{R}}$. If the task is node classification, $U_{L}=V_{L}$. If the task is link prediction, $U_{L}=\{u,v|(u,v)\in E_{train}\}$. By optimizing equation 6, we greedily add $f_{r_{\tau},u}$ to improve our final loss. Second-order approximation can be used to optimize equation 6 to quickly determine $w_{r_{\tau},0}$ and $w_{r_{\tau},1}$ for $f_{r_{\tau},u}$. The second-order approximation of equation 6 is:
$\mathcal{L}^{(\tau)}_{xg}\simeq$ $\sum_{u\in U_{L}}\sum_{v\in U_{L}}\Big{[}\ell\left(y_{u,v},\hat{y}^{(\tau-1)}_{u,v}\right)$
$+g_{u,v}f_{r_{\tau},u}(v)+\frac{1}{2}h_{u,v}f^{2}_{r_{\tau},u}(v)\Big{]}+\sum_{u\in U_{L}}\Omega(f_{r_{\tau},u}),$
where $g_{u,v}=\partial_{\hat{y}^{(\tau-1)}}\ell(y_{u,v},\hat{y}^{(\tau-1)})$ and $g_{u,v}=\partial^{2}_{\hat{y}^{(\tau-1)}}\ell(y_{u,v},\hat{y}^{(\tau-1)}).$ Since the Positive Sample Identification Problem is formulated as classification task, $\ell(\cdot,\cdot)$ here is the binary cross entropy, then $g_{u,v}=\sigma(\hat{y}^{(\tau-1)})-y_{u,v}$ and $h_{u,v}=\sigma(\hat{y}^{(\tau-1)})(1-\sigma(\hat{y}^{(\tau-1)})).$ $\ell\left(y_{u,v},\hat{y}^{(\tau-1)}_{u,v}\right)$ can be dropped because it does not affected by $w_{r_{\tau},0}$ and $w_{r_{\tau},1}$:
~$\widetilde{\mathcal{L}}^{(\tau)}_{xg}=$ $\sum_{u\in U_{L}}\sum_{v\in U_{L}}\Bigg{[}g_{u,v}f_{r_{\tau},u}(v)+\frac{1}{2}h_{u,v}f^{2}_{r_{\tau},u}(v)\Bigg{]}+\sum_{u\in U_{L}}\Omega(f_{r_{\tau},u})$
Let $I_{u,0}=\{v|s_{r_{\tau}}(u,v)\leq\eta_{r_{\tau},u}\wedge u,v\in U_{L}\}$ and $I_{u,1}=\{v|s_{r_{\tau}}(u,v)>\eta_{r_{\tau},u}\wedge u,v\in U_{L}\}$. The above equation can be re-written as
~$\widetilde{\mathcal{L}}^{(\tau)}_{xg}=\sum_{u}\Bigg{[}$ $\sum_{v\in I_{0}}g_{u,v}w_{r_{\tau},0}+\frac{1}{2}h_{u,v}w^{2}_{r_{\tau},0}+\sum_{v\in I_{1}}g_{u,v}w_{r_{\tau},1}$
$+\frac{1}{2}h_{u,v}w^{2}_{r_{\tau},1}+\frac{1}{2}\lambda(w^{2}_{r_{\tau},0}+w^{2}_{r_{\tau},1})\Bigg{]}$
The $w_{r_{\tau},0}$ and $w_{r_{\tau},1}$ that minimize ~$\widetilde{\mathcal{L}}^{(\tau)}_{xg}$ are
(7) $w_{r_{\tau},0}=\frac{-\sum_{u}\sum_{v\in I_{u,0}}g_{u,v}}{\sum_{u}(\sum_{v\in I_{u,0}}h_{u},v)+\lambda}$
(8) $w_{r_{\tau},1}=\frac{-\sum_{u}\sum_{v\in I_{u,1}}g_{u,v}}{\sum_{u}(\sum_{v\in I_{u,1}}h_{u},v)+\lambda},$
which are derived by solving $\partial_{w_{r_{\tau},0}}\widetilde{\mathcal{L}}^{(\tau)}_{xg}=0$ and $\partial_{w_{r_{\tau},1}}\widetilde{\mathcal{L}}^{(\tau)}_{xg}=0.$ Finally, the weights of regression tree $f_{r_{\tau},u}(v)$ for all query nodes is learned from equation 7 and equation 8. The details of how to train XGSampler iteratively is in Algorithm 1. Because the for loop at line 3 and line 9 are running through all pairs of nodes in the training node set $U_{L}$, the derivation of $w_{r,0}$ and $w_{r,1}$ can be implemented by the sparse matrix techniques in parallel. This makes the learning process of RelTree more efficient as discussed in Section 3.2. The ordered semantic relation set ¯=$\overline{{\mathcal{R}}}=\{r_{1},r_{2},\cdots,r_{|{\mathcal{R}}|}\}$ determines the training order of each regression tree. Instead of determining individual training order from $\{f_{r,u}|r\in{\mathcal{R}}\}$ for each $u$, a universal training order for all nodes is used. We determine training orders based on the probability ${\mathbb{P}}_{r}({\mathbb{I}}_{t}(u,v)=1|f_{r,u}(v)=1)$ In our experiment, we determine $\overline{{\mathcal{R}}}$ based on the probability descending order.
Finally, XGSampler can classify whether a node is the positive example for $u$ based on $f_{u}(v)$. $w_{r,0}$ and $w_{r,1}$ associated with each $f_{r,u}$ offer valuable insights into the significance of relation $r$ with respect to the target task. Higher values of $w_{r,0}$ and $w_{r,1}$ indicate greater influence in the determination of whether a node qualifies as a positive example. By deriving these weights for all semantic relations, XGSampler enhances the interpretability of node representations.
3.2. Computation Complexity Analysis
Because XGSampler is an additional process for selecting positive examples for Target-aware CL, we have to ensure that the extra cost is reasonable enough. A naive approach for sampling positive nodes via XGBoost for all $v\in{\mathcal{V}}$ can be $O(|{\mathcal{R}}||V|^{3}).$ Our learning approach on the contrary can be $O(|{\mathcal{R}}||V|^{2}\overline{B}),$ where $\overline{B}$ is the number of positive nodes. The number of parallel processes is set to be proportional to the number of nodes, i.e., $c|{\mathcal{V}}|$ where $c\in[0,1]$, the time complexity for XGSampler can be further reduced to $O(|{\mathcal{R}}||V|\overline{B}/c),$ in the graph. According to empirical computational results in Appendix C, we can learn node representations for a graph with half a million edges under a minute.
4. Semantic Relations
We describe the set of semantic relations ${\mathcal{R}}$ used in our experiments in following sections. We are interested in understanding the relations between two nodes $u\in{\mathcal{V}}$ and $v\in{\mathcal{V}}$. $N^{k}_{u}$ is the $k$-hop neighbors of $u$. Note that we use training data to calculate these similarity values. Below we specify the similarity measure for each semantic relation and commonly used definitions.
  • Adjacency matrix $A$ is a square matrix describing links between nodes. $A_{u,v}=1$ if $(u,v)\in\mathbb{E}$ else $A_{u,v}=0.$
  • $Y\in[0,1]^{|{\mathcal{V}}|\times C}$ is the one-hot encoding of class labels, where $C$ is the number of classes.
  • $X\in{\mathbb{R}}^{|{\mathcal{V}}|\times D}$ describes the attributes of all nodes.
Has Link (Link). (Kipf and Welling, 2017b)
(9) $s_{\text{link}}(u,v)=\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}[u,v],$
where $\widetilde{A}=A+I$ is the adjacency matrix $A$ with self-loop and $\widetilde{D}_{uu}=\sum_{v}\widetilde{A}_{uv}$ and $[u,v]$ indicates the $(u,v)$-entry of $\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}.$ This similarity is inspired by (Kipf and Welling, 2017b) in order to put weight on each link.
PageRank (PR). (Page et al., 1999) We can derive $\pi_{u}$ of personalized PageRank by solving the following equation
$\pi_{u}=\alpha P\pi_{u}+(1-\alpha)e_{u},$
where $P$ is the transition probability, $\alpha\in[0,1]$ and $e_{u}\in[0,1]^{|{\mathcal{V}}|}$ is a one-hot vector containing a one at the position corresponding $u$’s entry and zeros otherwise. The similarity in terms of PageRank is
(10) $s_{\text{PageRank}}(u,v)=\pi_{u,v}.$
Jaccard Similarity (Jacard Sim.). (Liben-Nowell and Kleinberg, 2007) Many studies have used this relation to predict links between node pairs. This relation uses the number of 1-hop common neighbors between two nodes to estimate the closeness between them.
(11) $s^{k}_{\text{jaccard}}(u,v)=\frac{|N^{k}_{u}\cap N^{k}_{v}|}{|N^{k}_{u}|\times|N^{k}_{v}|}.$
We use $k=1$ in our experiments.
Topology Similarity (Topology Sim.). (Dong et al., 2022) This similarity is similar to Jaccard Similary. Topology similarity represents the mutual information between the neighbors of two nodes. Please refer to (Dong et al., 2022) for mathematical definition of topology similarity.
Shortest Path Distance (Graph Distance). (Newman, 2003)
(12) $s_{\text{graph-distance}}(u,v)=\frac{\text{diameter}({\mathcal{G}})-d_{G}(u,v)+1}{\text{diameter}({\mathcal{G}})},$
where $d_{G}(u,v)$ is the shortest-path distance between $u$ and $v$. Note that $s_{\text{graph-distance}}(u,v)=\text{diameter}({\mathcal{G}})+1$ when $u$ and $v$ are disconnected in a graph, and hence $s_{\text{graph-distance}}(u,v)=0$.
Attribute Similarity. (Attr Sim.)
(13) $s^{k}_{\text{attribute similarity}}=\cos(x_{u},x_{v}),$
where $x_{u}\in{\mathbb{R}}^{D}$ is the node attributes of $u$ and $D$ is the dimension.
Attribute Distribution. (1-Hop Attr. Dist.) (Zhu and Ghahramani, 2002)
(14) $s^{k}_{\text{attribute distribution}}=\cos(x^{k}_{u},x^{k}_{v}),$
where $x^{k}_{u}=Normalize(A^{k}X)_{u}$ is the $k$-hop neighbor attribute distribution. We use $k=1$ in our experiments.
Label Distribution. (2-Hop Label Dist.) (Zhu and Ghahramani, 2002)
(15) $s^{k}_{\text{attribute distribution}}=\cos(y^{k}_{u},y^{k}_{v}),$
where $y^{k}_{u}=Normalize(A^{k}Y)_{u}$ is the $k$-hop neighbor label distribution. We use $k=2$ in our experiments.
Attribute Label Distribution. (Attr. Label Dist.)
(16) $s^{k}_{\text{attr. label distribution}}=\cos(\hat{y}^{k}_{u},\hat{y}^{k}_{v}),$
where $\hat{y}^{k}_{u}=Normalize(S^{k}_{\text{attr. dist.}}Y)_{u}$ is the label distribution of the $k$-hop attribute distribution similarity, $S^{k}_{\text{attr. dist.}}$. This similarity is especially useful for node classification to identify target-specific positive nodes.
Table 1. Accuracy of Node Classification Task. Statistically significant results are marked with * (p-value$<0.05$). The top scores are marked with bold blue, the 2nd scores are marked with blue, the 3rd scores are marked with violet.
Method Cora CiteSeer PubMed Computers Photo
DGI 84.42±0.17 69.37±2.1 84.35±0.08 83.59±0.2 91.92±0.06
SUGRL 82.86±0.42 72.59±0.19 85.17±0.16 85.28±1.11 92.85±0.09
ASP 84.58±0.35 73.53±0.01 84.0±0.01 84.03±0.01 90.52±0.01
PolyGCL 65.62±0.07 71.10±0.01 85.28±0.01 82.32±0.01 89.92±0.01
GCN 82.01±0.39 68.73±0.29 86.65±0.07* 88.01±0.18 92.59±0.1
GAT 83.01±0.36 69.16±0.31 85.08±0.11 88.16±0.12 93.0±0.08
DGCN 80.21±0.27 69.96±0.21 83.57±0.1 86.23±0.38 90.87±0.17
MetaPN 83.96±0.39 73.02±0.16 85.32±0.13 78.24±0.29 88.72±0.63
CoCoS 83.1±0.87 72.65±0.18 86.30±0.22 83.99±0.65 92.07±0.16
XTCL(GCN) 85.19±0.13* 73.55±0.14 86.09±0.09 89.54±0.09* 93.84±0.09*
XTCL(GAT) 84.13±0.24 73.26±0.32 85.47±0.08 88.50±0.10 93.15±0.12
5. Performance Evaluation
Our proposed model is Target-Aware CL with XGSampler, which we refer to as XTCL. Note that we can optimize any GNN method with XTCL. In this work, we demonstrate the effectiveness of our XTCL with GCN (Kipf and Welling, 2017b) and GAT (Velickovic et al., 2018). The dataset descriptions are in Table 3. In our empirical evaluation, we aim to answer:
  • RQ1. Does XTCL improve performance on node classification and link prediction compared to other SOTA methods?
  • RQ2. How supervision signals and training data size impact GNN performance?
  • RQ3. Is it important to make loss function task-aware?
  • RQ4. What semantic relations are important for each target task?
Dataset and Baseline We use widely used datasets to report the results of node classification and link prediction. Cora, CiteSeer and PubMed are citation networks (Yang et al., 2016) and Photo and Computers are product network (Shchur et al., 2018). If two products are often bought together, there is a link between them. The detailed of dataset statistics are shown in Table 3. For baseline methods, we compare our Target-aware CL with different training signals. DGI (Velickovic et al., 2019), SUGRL (Mo et al., 2022), ASP (Chen and Kou, 2023), and PolyGCL (Chen et al., 2024) are trained with their specified unsupervised or self-supervised signals. Variation of GNN models: GCN (Kipf and Welling, 2017b), GAT (Velickovic et al., 2018), DGCN (Cong et al., 2021), MetaPN (Ding et al., 2022), CoCoS (Xie et al., 2022), and SubGraph (Chamberlain et al., 2023) are trained with supervised signals based on their downstream tasks. See Appendix A to learn more about the detailed settings.
5.1. RQ1. Performance on Downstream Tasks
Node Classification
The performance results are shown in Table 1. XTCL(GCN) and XTCL(GAT) are GNNs trained with the proposed XTCL. XGSampler uses the available training labels to learn a model for each $u\in{\mathcal{V}}$, which we use to select positive examples for Target-aware CL. If node representations are generated in XTCL, self-supervised or unsupervised settings, Logistic Regression model is used as the final classifier for label assignment (with learned embeddings as input). We use $10\%$ labels for training and $90\%$ of labels for testing and the reported results are the average of five runs.
Table 1 shows that, with only small number of training labels ($10\%$) to train XGSampler, XTCL(GCN) significantly outperforms most of the state-of-the-art models in both unsupervised and (semi-)supervised settings. (See explanation for PubMed in Appendix D). The overall results in the table demonstrate that minimizing XTCL is similar to maximizing ${\mathcal{I}}({\bm{Y}};Z)$ and significantly improve the performance on node classification.
Table 2. AUC of Link Prediction. Statistically significant results are marked with * (p-value$<0.05$). The top scores are marked with bold blue, the 2nd scores are marked with blue, the 3rd scores are marked with violet.
Method Cora CiteSeer PubMed Computers Photo
DGI 93.36±0.2 94.17±0.21 94.18±0.09 84.0±4.58 93.03±0.08
SUGRL 93.74±0.16 96.89±0.09 95.81±0.06 91.89±0.26 95.50±0.06
ASP 91.45±0.01 95.30±0.01 93.47±0.01 89.02±0.01 91.19±0.02
PolyGCL 70.48±0.01 64.65±0.01 76.40±0.04 86.00±0.01 85.13±0.01
GCN 92.9±0.16 95.97±0.12 95.33±0.07 95.59±0.06 96.9±0.02
GAT 93.0±0.17 95.28±0.21 95.38±0.06 95.05±0.07 96.82±0.03
DGCN 91.56±0.16 89.46±4.45 95.36±0.11 93.73±0.05 95.64±0.42
SubGraph 87.01±1.33 88.95±2.20 95.08±0.15 98.00±0.05* 98.21±0.02
XTCL(GCN) 94.07±0.13* 97.13±0.06* 96.72±0.06* 97.32±0.03 98.38±0.08*
XTCL(GAT) 93.95±0.13 97.01±0.07 96.64±0.03 96.43±0.06 97.44±0.05
Link Prediction
For link prediction tasks, we compare to similar methods. We remove MetaPN and CoCoS because it is a semi-supervised model for node classification and cannot be adapted for link prediction. We follow the random split strategy used by (Chamberlain et al., 2023). To ensure the training data is limited, we use $60\%$ links for training and $40\%$ of links for testing and report the experiment results, averaged over five runs, in Table 2. The observations from Table 2 is similar to that of Table 1. By training GCN with XTCL, its performance can be improved significantly also for link prediction. This indicates that, by increasing the probability to select proper positive examples for link prediction, Target-aware CL successfully to improve performance over other methods.
5.2. RQ2. Comparing Various Supervision Signals
In this experiment, we explore whether XGSampler can improve XTCL performance on node classification as the training label size increases. The result is shown in Figure 2. We use the same $20\%$ testing labels for all evaluated models, and the results are the average of five runs. We train GCN with various objective functions, and Logistic Regression is used as the classifier for all CL models. The learning objectives of Ceiling Supervised Contrastive Loss (Ceiling SCL), Naive Target-Aware CL (Naive TCL) are in Def. 5.1 and Def. 5.2.
Figure 2 shows that XTCL outperforms Naive TCL because XGSampler increases the probability that ${\mathbb{I}}_{t}(u,v)=1$ for a positive example $v$. However, as the training size increases, the performance gap between XTCL and Naive TCL becomes smaller. This is because Naive TCL has access to more training labels, allowing it to sample more positive nodes that ${\mathbb{I}}_{t}(u,v)=1$ as positive examples. A similar observation can be applied to SCL. Furthermore, XTCL can sometimes be comparable to the Ceiling SCL because XTCL enhance model generalization further by incorporating abandon graph signals during node representation learning.
Definition 0.
Ceiling Supervised Contrastive Loss (Ceiling SCL). We use Ceiling SCL to refer to a method that uses ground truth labels to identify positive examples, ie.
$\left\{v_{p_{b}}\right\}_{b=1}^{B}\subseteq\{v\>|\>{\mathbb{I}}_{t}(u,v)=1,v\in V\}$
The difference between SCL in Def. 2.2 and Ceiling SCL is that Ceiling SCL can utilize the label of $y_{u}$ even when $u$ is not in $V_{L}$.
Definition 0.
Naive Target-Aware CL (Naive TCL). Naive TCL is a combination of SSCL in Def. 2.1 and SCL in Def. 2.2. It uses Def. 2.1 to select positive examples when $u\notin V_{L}$ If $u,v\in V_{L}$, the positive examples are sampled from the set $\{v\>|\>{\mathbb{I}}_{t}(u,v)=1,u,v\in V_{L}\}$.
5.3. RQ3. Importance of Task-Aware Loss Function
In Section 3.1, we briefly mentioned that the selection of the loss function will impact the model generalization for a target task. Our main argument is that minimizing the value of a loss function is not equivalent to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. Consequently, model performance and generalization degrades severely.
Figure 2. Learning curves of GCN trained with various supervision signals. Our XGSampler is effective because XTCL continues to outperform others when the training size increases.
(a) Node Classification ACC.
(b) Train ACC - Test ACC.
Figure 3. Results on semi-synthetic data when varying the dependency between node labels and a graph signal (1-Hop Attr. Dist.). The results show that minimizing losses, which are not task-aware, is not equivalent to maximizing the mutual information ${\mathcal{I}}({\bm{Y}};Z)$, which decreases model generalization on node classification.
To empirically investigate the above conjecture, we conducted an experiment using semi-synthetic data generated from Cora. The target task is classification task. First, we choose a semantic relation $\hat{r}$ subject to that knowing $s_{\hat{r}}(u,v)$ does not inform the probability of $y_{u}=y_{v}$ and ${\mathcal{I}}(S_{\hat{r}};Z)$ is less affected by optimizing the loss function that generates $Z$. Then, we gradually perturb some node labels in ${\bm{Y}}$ such that ${\mathcal{I}}({\bm{Y}}^{\prime};S_{\hat{r}})>{\mathcal{I}}({\bm{Y}};S_{\hat{r}})$. ${\bm{Y}}^{\prime}$ is the new node label set with some labels perturbed. This will create a situation that minimizing the value of a loss function is not equivalent to maximizing ${\mathcal{I}}({\bm{Y}}^{\prime};Z)$. Because ${\mathcal{I}}(S_{\hat{r}};Z)$ stays similar by optimizing the loss function, ${\mathcal{I}}({\bm{Y}}^{\prime};Z)$ drops when ${\mathcal{I}}({\bm{Y}}^{\prime};S_{\hat{r}})$ increases. We want to observe how this situation affect the model performance. When the number of perturbed nodes increase, ${\mathcal{I}}({\bm{Y}}^{\prime};S_{\hat{r}})$ will increases accordingly in Figure 3. The $\hat{r}$ for this experiment is 1-Hop Attr Dist. The detail of label perturbation process and the experiment settings of Figure 3 is in Appendix B.
We compare GNN optimized by these loss functions: self-supervised (SUGRL (Mo et al., 2022)), supervised (GCN (Kipf and Welling, 2017b)), semi-supervised (MetaPN (Ding et al., 2022)), and our target-aware contrastive loss (XTCL). The goal is to observe how the performance of different models are affected when ${\mathcal{I}}({\bm{Y}}^{\prime};S_{\hat{r}})$ (number of perturbed nodes) increases.
Figure 4. Importance Weights of Semantic Relations. This figure demonstrates the importance of these graph signals (x-axis) to each target task and how our XGSampler adapt XTCL to improve model performance.
Both self-supervised methods (SUGRL) and semi-supervised methods (MetaPN) have greater performance degradation. Because the generated representation does not capture 1-Hop Attr Dist by maximizing their loss functions. in the embedding space. This illustrate that, since maximizing ${\mathcal{I}}(S_{r};Z)$ does not maximize ${\mathcal{I}}({\bm{Y}}^{\prime};Z)$ simultaneously, the performance drops severely. As supervised methods (GCN), its performance is subject to the data distributions in training and testing, which is the overlapping area between two red circles in Figure 1. GCN in Figure 3 overfits to training data because the number of training labels is insufficient (10% of the data). The severity of GCN’s overfitting is shown in Figure 3(b).
In contrast, our XTCL shows relatively robust performance in Figure 3. Because XTCL can actively adapt the number of positive examples for each $r\in{\mathcal{R}}_{f}$ w.r.t the target task, minimizing XTCL is now similar to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. Therefore, the generated node representations can generalize better w.r.t target task and largely reduced the chance to overfitting to training data.
5.4. Importance Weights of Semantic Relations
The weights $w_{r,0}$ and $w_{r,1}$ learned from the XGSampler indicate the importance of semantic relation $r$ in determining whether a node is a positive example for $u$. This weights also show that the importance of each $r$ to a target task. The importance weight in Figure 4 is $\max(|w_{r,0}|,|w_{r,1}|)$ for each relation. The definition of each semantic relation is in Section 4. Figure 4 demonstrates that different tasks (e.g. node classification vs. link prediction) rely on different semantic relations to determine positive examples, which means one semantic relation which is important to a target task may not be as critical as another target task. This result corresponds to our main argument that uniformly maximizing ${\mathcal{I}}(S_{r};Z)~{}\forall r\in{\mathcal{R}}_{f}$ is not always equivalent to maximizing ${\mathcal{I}}({\bm{Y}};Z)$. Figure 4 also demonstrate that XGSampler improves the interpretability of the learned embeddings, by illustrating which semantic relation is critical for a target task.
6. Related Work
Contrastive Learning have been applied to learning representations for various type of data. The selection of appropriate positive and negative examples is a crucial step in optimizing contrastive loss. In the context of Computer Vision (CV), various image augmentation techniques such as rotation and clipping have been proposed to generate positive examples from the query image (Le-Khac et al., 2020; Chen et al., 2020). Conversely, in the field of Neural Language Processing (NLP), several pre-trained models are employed to generate positive samples for words or sentences  (Gao et al., 2021). These data augmentation approaches have demonstrated significant improvements in tasks related to CV and NLP.
Numerous task-aware contrastive objectives have been proposed for word and image representations. The initial work is Supervised Contrastive Learning  (Khosla et al., 2020), which selects positive and negative examples based on the available labels in the training data. The performance of text  (Suresh and Ong, 2021; Grover et al., 2022) and image classification tasks (Khosla et al., 2020) has been enhanced through the use of supervised contrastive loss. The SCL depicted in Figure 3 and Figure 2 adopts a similar setting to supervised contrastive learning, and both figures illustrate that the performance is constrained by unlabeled nodes due to the absence of an appropriate method for generating positive samples for those nodes.
Several graph augmentation techniques, e.g. adding/dropping edges (Suresh et al., 2021b; Sun et al., 2020)or modifying attributes (Velickovic et al., 2019; You et al., 2020), have been proposed. However, these graph augmentation techniques are specifically designed for entire graphs, whereas a single node is more sensitive to local changes made to the graph through these techniques. Consequently, many methods are proposed to sample positive examples based on structural or topological similarity in the graphs (Mo et al., 2022; Dong et al., 2022). However, the relationship between their unsupervised contrastive signals and the objective of a target task remains unclear, which may lead to arbitrary performance deterioration, as demonstrated in Table 1 and Table 2.
7. Conclusion
To ensure that minimizing self-supervised losses is similar to maximizing the mutual information between a target task and node representations, we propose Target-Aware Contrastive Loss with XGBoost Sampler (XTCL). The proposed XGSampler increases the probability of selecting nodes that enhance the mutual information between node representation and a target task as positive examples for contrastive loss. It learns how the target task relates to the semantic relations in the graph and actively adjusts the number of positive examples for each semantic relation. The performance results on both node classification and link prediction have shown that XTCL can significantly outperform the state-of-the-art unsupervised and self-supervised models and significantly outperform supervised models in $90\%$ of our experiments. Additionally, the interpretability of XTCL is enhanced by the weights assigned to each semantic relation. These weights indicate the importance of each semantic relation to a target task.
Appendix A Experiment Setup
All codes and scripts are run in the environment with Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and NVIDIA A10 GPU. The scripts are available on GitHub [^2] Github link will be provided after publication to preserve anonymity.. We refer to the official repository to implement all the baseline methods. We tuned for best parameter settings using grid search and the parameter settings for our models is in our GitHub repo and the dataset statistics are in Table 3.
Table 3. Data Statistics. Cora, CiteSeer and PubMed are from (Yang et al., 2016). Computers and Photo are from (Shchur et al., 2018).
Dataset Node Edge Node Class Node Attr.
Cora 2708 10556 7 1433
CiteSeer 3327 9104 6 3703
PubMed 19717 88648 3 500
Photo 7650 238162 8 745
Computers 13752 491722 10 767
Appendix B Label Perturbation Process
We gradually increase the influence of $\hat{r}$ on the classification task by determining a portion of node labels using $\hat{r}$. Initially, when the level of perturbation is low, $\hat{r}$ is less relevant to the node labels, i.e. ${\mathcal{I}}({\bm{Y}};S_{\hat{r}})$ is low. We gradually increase the influence of $\hat{r}$ on node labels by setting the new label of node $u$ the same as $v$ if $s_{\hat{r}}(u,v)>\eta_{\hat{r}}$ for a pair of node $u$ and $v$. As the number of perturbed pairs increases, ${\mathcal{I}}({\bm{Y}}^{\prime};S_{\hat{r}})$ increases. By varying the number of perturbed nodes, we can control ${\mathcal{I}}({\bm{Y}}^{\prime};S_{\hat{r}})$.
To increase the influence of semantic relation $\hat{r}$ on node labels, the label perturbation must happen in a collective manner, e.g. changing the labels of nodes which have high similarity based on the selected semantic relation $\hat{r}$ together. The key lies in that the graph induced by the similarity matrix of $\hat{r}$ should be homophilc to the perturbed nodes (Suresh et al., 2021a; Klicpera et al., 2019; Cong et al., 2021) to enhance the dependency between $r$ and node labels. We select one semantic relations defined in Section 4 as the semantic relation to perturb node labels. Our perturbation process is described in Algorithm 2.
The results in Figure 3 is based on the semi-synthetic data generated from Cora (see Table 3). We keep the structure of Cora and only determine the labels based on 1-hop Attr. Dist., which is less relevant to ${\bm{Y}}$ (also shown in Figure 4) and to the following loss functions. For each point in Figure 3, we generate three Cora graphs with different perturbed labels ${\bm{Y}}^{\prime}$ based on $\hat{r}$, and we create five train/test splits (0.1 for training/0.9 for testing) for each graph with labels perturbed. Given one perturbed graph with one train/test split, we train all models based on the training data and evaluate them with test data. The accuracy (ACC) quantifies whether the predicted label for a test node is the same as its label in the test data. The reported accuracy and standard errors are derived from these $15$ runs.
Algorithm 2 Label Perturbation Process
0:  the selected signals $\hat{r}\in\mathbb{R}$, its similarity function $s_{\hat{r}}$, similarity threshold $\{\eta_{\hat{r},u}|u\in{\mathcal{V}}\}$, number of perturbation nodes $p$, collective ratio $q$, original labels ${\bm{Y}}$
0:  perturbed node labels ${\bm{Y}}^{\prime}$
1:  ${\bm{Y}}^{\prime}={\bm{Y}}$
2:  Initialize stack $U$
3:  Randomly sample $u\in{\mathcal{V}}$ and put it into ${\mathcal{V}}^{\prime}$
4:  while $|{\mathcal{V}}^{\prime}|<p$ do
5:     Pop $u$ from $U$
6:     if $u\notin{\mathcal{V}}^{\prime}$ then
7:        for all $\{v|s_{\hat{r}}(u,v)>\eta_{\hat{r},u}\}$ do
8:           $q^{\prime}\sim\text{Uniform}(0,1)$
9:           if $q^{\prime}<=q$ then
10:              $y^{\prime}_{v}\leftarrow y^{\prime}_{u}$
11:           end if
12:        end for
13:        ${\mathcal{V}}^{\prime}={\mathcal{V}}^{\prime}\cup u$
14:        $U=U\cup\{v|s_{\hat{r}}(u,v)>\eta_{\hat{r},u}\}$
15:     end if
16:  end while
Appendix C Empirical Computation Time Analysis
Table 4. Computational Time in seconds.
Cora CiteSeer PubMed Photo Computers
Naive XGBoost 4.05 4.89 66.19 17.76 41.81
Our XGSampler 0.77 0.9 10.59 2.31 5.65
GCN Training 4.15 5.95 40.79 31.33 34.40
According to the results in Table 4, the average computational time for learning XGSampler per node is reduced by $84\%$, compared to Naive XGBoost with time complexity $O(|{\mathcal{R}}||V|^{3}).$ The total time for learning only increases by an average of $17\%$ when comparing the GCN training time and the preprocessing time for XGSampler. After setting the learning processes in parallel, the computational time to learn XGSampler is linear to number of nodes in a graph.
Appendix D PubMed Performance Discussion
XTCL(GCN) achieves slightly lower performance compared to GCN in PubMed, but we conjecture that this is due to the smaller number of class label values which enables lower generalization error for GCN.
We conducted additional experiments to test this hypothesis in Section 5.1 We attempted to decrease the number of classes for training and prediction on Cora and observed how the performance varied with the number of classes. The averages of 5 testing runs are as follows:
Num Classes Accuracy
3 91.7
4 86.7
5 85.8
6 83.7
7 82.0
As the number of classes decreases, the performance of GCN increases accordingly. This is an initial result to test the hypothesis, and a comprehensive investigation is required to confirm it.